OpenStudy (anonymous):

prove the identity that (tan^2 x) - (sin^2 x) = (tan^2 x)(sin^2 x)

sam (.sam.):

Replace with an equivalent expression using the fundamental identities. ((sin^(2)x)/(cos^(2)x))-(sin^(2)x) Multiply -1 by each term inside the parentheses. (sin^(2)x)/(cos^(2)x)-sin^(2)x Combine (sinx^(2))/(cosx^(2))-sinx^(2) into a single expression by finding the least common denominator (LCD). The LCD of (sinx^(2))/(cosx^(2))-sinx^(2) is cosx^(2). (sin^(2)x-sin^(2)xcos^(2)x)/(cos^(2)x) factor out the GCF of sinx^(2) from sinx^(2)-sinx^(2)cosx^(2). (sin^(2)x(1-cos^(2)x))/(cos^(2)x) Multiply sinx^(2)*sinx^(2) to get sinx^(4) in the numerator. (sin^(4)x)/(cos^(2)x) multiply -1 by each term inside the parentheses. (sin^(2)x)/(cos^(2)x)-sin^(2)x Combine (sinx^(2))/(cosx^(2))-sinx^(2) into a single expression by finding the least common denominator (LCD). The LCD of (sinx^(2))/(cosx^(2))-sinx^(2) is cosx^(2). (=sin^(2)x-sin^(2)xcos^(2)x)/(cos^(2)x) factor out the GCF of sinx^(2) from sinx^(2)-sinx^(2)cosx^(2). =(sin^(2)x(1-cos^(2)x))/(cos^(2)x) Multiply sinx^(2)*sinx^(2) to get sinx^(4) in the numerator. =(sin^(4)x)/(cos^(2)x) rewrite the expression using the (sinx)/(cosx)=tanx identity. =sin^(2)xtan^(2)x

sam (.sam.):

quite long, lol

OpenStudy (anonymous):

whoh ok thanks

OpenStudy (experimentx):

(sin(x))^2/(cos(x))^2 - (sin(x))^2 =>(sin(x))^2((1/(cos(x))^2)-1) 1/(cos(x))^2) = sec^2x as we know (sec(x))^2-1) = (tan(x))^2 hence we have LHS

OpenStudy (anonymous):

gracias por ayudandome